When you have two unequal denominators in fractions, how do you add the two together without increasing them? Would that even be possible? Please help.
Found 2 solutions by MathLover1, Alan3354:
Answer by MathLover1(20850)


If the denominators are not the same, then you have to use equivalent fractions which do have a common denominator . To do this, you need to find the least common multiple (LCM) of the two denominators. To add fractions with unlike denominators, rename the fractions with a common denominator. Then add and simplify.
how do you add the two together without increasing them?- if the denominators are not the same, it's not possible
only if the denominators are the same, you can add the two together without increasing them

Answer by Alan3354(69443)

When you have two unequal denominators in fractions, how do you add the two together without increasing them? Would that even be possible?
-----------
Depends on the DENs.
eg,
1/3 + 4/6 = 1/3 + 2/3 = 1 --- no increase of the DENs